OpenStudy (vivianshock1):

The area of a rectangle is 56 square centimeters. The length of the rectangle is 8 centimeters, and the width is three less than a number. What is the number?

OpenStudy (slenderman):

(56/8)=x-3 56/8= 7= 10-3

OpenStudy (zpupster):

Area= length*width plugging in 56= 8(n-3)

OpenStudy (akshat):

let number be x then width = x-3 also area of sq = area rectangle 56=8(x-3) 56/8=x-3 7=x-3 so x= 10

OpenStudy (welshfella):

@Akshat - plz dont give the answer - gOS adviises us to guide the use to the answer